how do i solve this word problem ? .

The area of a rectangle is 24 square meters. Find the length and width of the rectangle if its length is 2 meters greater than its width. Use an equation and the formula ® area of a rectangle = (width)(length).

Found 2 solutions by stanbon, checkley71:
Answer by stanbon(48558)

The area of a rectangle is 24 square meters. Find the length and width of the rectangle if its length is 2 meters greater than its width. Use an equation and the formula ® area of a rectangle = (width)(length).
------------------------
Let width be "x" meters
Then length is "x+2" meters
EQUATION:
x(x+2) = 24
x^2+2x-24 = 0
(x+6)(x-4)= 0
x = 4 meters (width)
x+2 = 6 meters (length)
==============================
Cheers,
Stan H.

Answer by checkley71(8403)

L=W+2
24=2(W+2)+2W
24=2W+4+2W
4W=24-4
4W=20
W=20/4
W=5 FOR THE WIDTH.
L=5+2=7 FOR THE LENGTH.
PROOF
24=2*7+2*5
24=14+10
24=24